Sarah and her sister were born 12 years apart. If sarah is 3 times older than her sister, how old are the girls?
Answer by Maths68(1474)

Let
Present age of Sarah = s
Present age of sister = t
Sarah and her sister were born 12 years apart.
s-t=12
s=12+t.........(1)
sarah is 3 times older than her sister
s=3t...........(2)
Put the above value of s in (1)
3t=12+t
3t-t=12
2t=12
2t/2=12/2
t=6
Put the value of t in (2)
s=3t
s=3*6
s=18
Present age of Sarah = s = 18 years old
Present age of sister = t = 6 years old